# Think in Safety: Unveiling and Mitigating Safety Alignment Collapse in Multimodal Large Reasoning Model

### v2

@@ Abstract
The rapid development of Multimodal Large Reasoning Models (MLRMs) has demonstrated broad application potential, yet their safety and reliability remain critical concerns that require systematic exploration. To address this gap, we conduct a comprehensive and systematic safety evaluation of 11 MLRMs across 5 benchmarks and unveil prevalent safety degradation phenomena in most advanced models. Moreover, our analysis reveals distinct safety patterns across different benchmarks: significant safety degradation is observed across jailbreak robustness benchmarks, whereas safety-awareness benchmarks demonstrate less pronounced degradation. In particular, the long thought process in some scenarios even enhances safety performance. Therefore, it is a potential approach to address safety issues in MLRMs by leveraging the intrinsic reasoning capabilities of the model to detect unsafe intent. To operationalize this insight, we construct a multimodal tuning dataset that incorporates a safety-oriented thought process. Experimental results from fine-tuning existing MLRMs with this dataset effectively enhances the safety on both jailbreak robustness and safety-awareness benchmarks. This study provides a new perspective for developing safe MLRMs. Our dataset is available at https://github.com/xinyuelou/Think-in-Safety.
@@
Xinyue Lou1111Equal contribution. Contact e-mail: louxinyue@bjtu.edu .cn, youlee@bjtu.edu.cn, You Li1111Equal contribution. Contact e-mail: louxinyue@bjtu.edu .cn, youlee@bjtu.edu.cn, Jinan Xu1, Xiangyu Shi1, Chi Chen2, Kaiyu Huang1 1 Beijing Jiaotong University, Beijing, China 2 Tsinghua University, Beijing, China
22footnotetext: Corresponding author: kyhuang@bjtu.edu.cn
@@ 1 Introduction
RQ1: How does MLRMs affect safety on different types of benchmarks compared with Multimodal Large Language Models (MLLMs)?
@@
Gemini2.0-Flash-Thinking 89.80 73.48 50.36 39.16 57.32
Claude3.7-Sonnet-Thinking 32.20 33.94 0.49 59.24 57.76
QVQ-Preview 70.80 69.29 37.86 34.13 50.68
Skywork-R1V 85.80 72.68 35.28 35.33 50.42
Llama 3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
LlamaV-o1 59.40(+3.60) 53.93(+15.48) 13.57(+7.86) 33.93(-3.20) 51.59(-0.67)
Llama 3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
LLaVA-CoT 84.80(+29.00) 72.26(+33.81) 33.57(+27.86) 26.95(-10.18) 51.67(-0.59)
Llama 3.2-vision-11B(base) 55.80 38.45 5.71 37.13 52.26
Mulberry-Llama 67.40(+11.60) 64.70(+26.25) 13.21(+7.50) 37.72(+0.59) 54.09(+1.83)
Qwen2.5-VL-3B(base) 66.27 66.18 12.14 24.55 52.35
LMM-R1 69.80(+3.53) 68.15(+1.97) 18.21(+6.07) 21.56(-2.99) 53.02(+0.67)
Qwen2.5-VL-7B(base) 67.20 66.49 12.50 29.94 50.02
R1-Onevision 72.20(+5.00) 79.57(+13.08) 32.14(+19.64) 17.31(-12.63) 48.94(-1.08)
InternVL-2.5-8B(base) 71.40 59.64 15.00 28.14 50.84
$M$-Eureka 72.20(+0.80) 60.12(+0.48) 11.79(-3.21) 28.14(-0.00) 50.59(-0.25)
Kimi-VL(base) 80.40 47.74 22.50 25.00 50.44
Kimi-VL-Thinking 87.00(+6.60) 61.49(+13.75) 33.93(+11.43) 35.93(+10.93) 51.42(+0.98)
@@ 2.1 Evaluation Settings
In safety-awareness tasks, models need to jointly reason over both visual and textual inputs to infer user intent, identify potential safety risks, and assess whether the input should be treated as safe or unsafe. These tasks pose significant challenges to the multimodal reasoning and safety alignment capabilities of models. In contrast, jailbreak robustness benchmarks involve adversarial attacks, such as the inclusion of maliciously crafted prompt, aimed at circumventing the safety constraints of models. Given that these two task categories examine safety alignment from fundamentally different perspectives, we analyze these results separately. Further details regarding the datasets are provided in Appendix A.1.
@@
We evaluate a total of 11 MLRMs, including both proprietary models and open-source models along with their corresponding base models, such as Kimi-VL-Thinking Team et al. (2025), R1-Onevision Yang et al. (2025), Mulberry Yao et al. (2024), and LlamaV-o1 Thawakar et al. (2025). Detailed information of the models are provided in Appendix A.3. For reasoning models that have undergone extensive fine-tuning—via either supervised learning or reinforcement learning—we additionally evaluate their corresponding base models. This approach allows us to trace safety-related changes across model variants. All experiments are conducted using four NVIDIA A100-80G GPUs.
@@
For jailbreak robustness benchmarks, we adopt the standard Attack Success Rate (ASR) metric, which quantifies the percentage of instances where the model produces harmful outputs with its safety mechanisms circumvented. A lower ASR indicates stronger safety alignment. For safety-awareness benchmarks, we follow established evaluation protocols to compute a safety score, where higher scores denote better safety performance. Further evaluation settings are detailed in Appendix A.1. Following the settings of MM-SafetyBench, MSS, and SIUO utilizing OpenAI API, we employ GPT-4o-mini Hurst et al. (2024) as the evaluator and adopt identical configurations to ensure consistency across assessments.
@@ 2.2 Safety Results and Inspection
Finding 1 ($\Rightarrow$ RQ1): The safety performance of MLRMs requires improvement and exhibits a notable security degradation phenomenon.
@@
As shown in Table 1, existing multimodal reasoning models exhibit significant security vulnerabilities, with most open-source models performing unsatisfactorily across various benchmarks. In particular, Kimi-VL-Thinking, Skywork-R1V and LLaVA-CoT successfully defend against only about 15% of malicious queries on Figstep. This finding highlights the necessity of ensuring safety alignment alongside improvements in multimodal reasoning capabilities.
Security Degradation.
Most reasoning models exhibit a notable increase in ASR on jailbreak robustness benchmarks relative to their corresponding foundation models, with an average increase of approximately 11.53%. This suggests that the internal safety alignment mechanisms are compromised during the process of enhancing reasoning capabilities, thereby rendering the models more susceptible to jailbreak attacks. In contrast, the degradation of safety performance is less evident in safety-awareness benchmarks. In some cases, models even demonstrate improved safety performance after engaging in extended thought process. A detailed analysis of this phenomenon is presented in Finding 2.
@@
Performance on Safety-Awareness.
As shown in Table 1, the safety score of models such as Kimi-VL-Thinking Team et al. (2025), LMM-R1 Peng et al. (2025b), and MM-Eureka Meng et al. (2025) are higher than their corresponding base models on safety-awareness benchmark. The effect is particularly pronounced for Kimi-VL-Thinking model, which achieves a 10.9% improvement on the SIUO benchmark. As illustrated in Figure 2, there is a difference in focusing center according to the responses generated by these models and their corresponding base models. Reasoning models, when engaging in long reasoning, tend to focus more on potentially unsafe details present in both the visual and textual modalities. These models improve the safety performance through internal self-reflection and error-correction mechanisms. This encouraging finding offers a promising direction for mitigating safety degradation by leveraging the innate reasoning capabilities of the models themselves.
@@
Performance of Thought Process.
Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated even if the thought process is secure.
@@
As shown in Table 3, the results reveal that the potential risk arising from the thought process does not always show up consistently in the final answer. For example, the safe thought process leads to the unsafe answer in QVQ-Preview, with respect to 46.60% proportion. In specific examples, when the model meets with unsafe intents, it is not engaged in in-depth thinking to generate unsafe content in thought while directly generate unsafe answers in response. The inconsistency between safe thought and unsafe answer is due to the potential security risks inherent in the model. On the contrary, the unsafe thought process also results in the safe final answer. It is attributed to the internal self-reflection and error-correction mechanisms of MLRMs, which enables models to eventually recognize the harmful nature of the intended prompt. The comparative results also provide potential insights for mitigating the issue of safety degradation in MLRMs.
@@ 3 Data Construction
Although existing studies have introduced several datasets for multimodal safety alignment, most of them Zong et al. (2024); Ding et al. (2025b) consist of brief responses that lack explicit thought process. The fine-tuning method based on these datasets struggles with retaining the inherent reasoning-chain advantages of MLRMs. Furthermore, arriving at conclusions without engaging in a safety-oriented thought process weakens the effectiveness and robustness of safety defense mechanisms. To address these issues, we propose a method for building a multimodal alignment dataset named TiS (Think in Safety), distinguished by its safety-oriented thought process and various safety-related topics. Through this method, we aim to leverage the innate reasoning capabilities of MLRM to improve its safety alignment.
As shown in Figure 5, we employ a multi-stage pipeline to construct our safety alignment dataset TiS. We begin by collecting safety-related topics and generating image captions, then explicitly incorporating long CoT reasoning into question answering. After a filtering procedure, we finally obtain the dataset. To the best of our knowledge, TiS is the first safety dataset with the ability to retain reasoning chain for MLRMs. We provide a detailed illustration of the entire pipeline in Appendix B.1.
@@
Due to the limitation of existing multimodal safety alignment datasets, we aim to construct a dataset enriched with safety-oriented thought process. However, considering the inherent security vulnerabilities of existing models and the resource constraints, constructing such a dataset entirely from scratch is largely impractical. Therefore, we augment existing datasets with short responses by adding structured and security-oriented thought process. Specifically, we select two multimodal safety alignment datasets as our original data sources: VLGuard Zong et al. (2024) and SPA-VL Zhang et al. (2024). Rather than directly utilizing the entire datasets, we retain only the high-quality instances and ensure a balanced distribution across diverse safety-related topics.
@@
To obtain a reliable thought process, we utilize the text-only reasoning model, Deepseek-R1 Guo et al. (2025), because instruction models often struggle to produce coherent reasoning, while MLRMs tend to suffer from alignment collapse under various conditions. Meanwhile, close-source proprietary models are equipped with defensive safety guardrail, which hinders access to their internal reasoning processes particularly for safety-related questions. To address the modality gap inherent in this text-only models, the visual content is converted into detailed image captions using Qwen2.5-VL-72B Bai et al. (2025).
@@
To ensure the thought process generated by LRMs is aligned with human values, we explicitly add safety guidelines in the generation process, inspired by the concept of deliberative alignment Guan et al. (2024). Specific safety standards for various domains (law, medical, etc) are also clarified. Previous methods often require models to infer implicit safety rules from large volumes of training examples, which suffers from low efficiency and limited generalization. By contrast, our method provides sufficient safety standards for safety-centric reasoning content generation, facilitating deeper safety alignment and understanding rather than the first few tokens Qi et al. (2024). More details and prompt used in this step are provided in Appendix B.3.
@@
After obtaining the safety-oriented reasoning processes, we perform an additional modification and filtering step to ensure the quality and correctness of the data, making them better suited for safety enhancement. This step addresses issues such as inaccuracies (e.g., responses that explicitly state “according to the caption”) and misalignments in reasoning. We collaboratively employ Deepseek-V3 Liu et al. (2024a) alongside human annotators to assess data quality and filter inappropriate instances.
@@ 4.2 Results
As shown in Table 4, both R1-Onevision and LLaVA-CoT demonstrate improved safety alignment fine-tuned on TiS, substantially outperforming prior datasets. Specifically, our dataset TiS enhances safety performance on Figstep and SIUO by 30% compared to the best alternative baseline, effectively enabling the models to leverage their reasoning capabilities for deeper analysis and unsafe intention detection. Furthermore, since our dataset incorporates thought process that closely align with the data distribution used in MLRM training, the fine-tuned model preserved the original capability to generate coherent reasoning pathways. More case studies are provided in Appendix D.4.
@@ 4.3 Discussion
To investigate whether the proposed thought process can further enhance the safety performance of MLRMs, we conduct experiments by removing the thought component from TiS dataset, retaining only the answer portion. As shown in Table 4, retaining the thought process leads to more substantial improvements in safety performance compared to using answer-only data, with the exception of MM-SafetyBench. For certain instances of MM-SafetyBench, the model’s deeper reasoning leads it to categorize the input as neutral, which leads to the increase of ASR.The answer-only data also exhibits favorable performance, suggesting that the construction of more comprehensive and higher-quality datasets facilitates the safe fine-tuning of MLRMs. However, this can lead to generated responses that not only lack the thought process but also consist solely of brief replies such as “I’m sorry, I can’t assist it.”.More case studies are provided in Appendix D.3. Furthermore, this also demonstrates the feasibility of leveraging the reasoning capabilities of MLRMs to enhance the alignment with safety objectives.
@@
To further assess the impact of Tis on enhancing the security of internal reasoning, we separately evaluate the safety of both thought process and final answer, as shown in Table 5. The results demonstrate the effectiveness of TiS in safeguarding both thought & answer response, with a substantial reduction in the ASR across all categories. Notably, TiS effectively reduces the proportion of cases in which both thought process and final answer are unsafe, decreasing from 42.80% to 0.20% on R1-Onevision, and from 54.60% to 0.40% on LLaVA-CoT. Furthermore, in scenarios where unsafe content appears solely in the intermediate thinking, our fine-tuning approach also achieves a marked improvement, reducing such instances from 14.20% to 2.60%. These results indicate that TiS largely enhances the reliability and security of the model’s internal reasoning process.

### v4

@@ Abstract
The rapid development of Multimodal Large Reasoning Models (MLRMs) has demonstrated broad application potential, yet their safety and reliability remain critical concerns that require systematic exploration. To address this gap, we conduct a comprehensive and systematic safety evaluation of 13 MLRMs across 5 benchmarks and unveil prevalent safety degradation phenomena in most advanced models. Moreover, our analysis reveals distinct safety patterns across different benchmarks: significant safety degradation is observed across jailbreak robustness benchmarks, whereas safety-awareness benchmarks demonstrate less pronounced degradation. In particular, the long thought process in some scenarios even enhances safety performance. Therefore, it is a potential approach to address safety issues in MLRMs by leveraging the intrinsic reasoning capabilities of the model to detect unsafe intent. To operationalize this insight, we construct a multimodal tuning dataset that incorporates a safety-oriented thought process. Experimental results from fine-tuning existing MLRMs with this dataset effectively enhance the safety on both jailbreak robustness and safety-awareness benchmarks. This study provides a new perspective for developing safe MLRMs. [^1] Our dataset is available at https://github.com/xinyuelou/Think-in-Safety.
@@
Xinyue Lou1,2, You Li1,2, Jinan Xu1,2, Xiangyu Shi1,2, Chi Chen3, Kaiyu Huang1,2222Kaiyu Huang is the corresponding author. 1Key Laboratory of Big Data & Artificial Intelligence in Transportation (Beijing Jiaotong University), Ministry of Education 2School of Computer Science and Technology, Beijing Jiaotong University 3Tsinghua University {louxinyue,kyhuang}@bjtu.edu.cn
@@ 1 Introduction
RQ1: How do MLRMs affect safety on different types of benchmarks compared with Multimodal Large Language Models (MLLMs)?
@@
[…]
@@ 2.1 Evaluation Settings
In safety-awareness tasks, models need to jointly reason over both visual and textual inputs to infer user intent, identify potential safety risks, and assess whether the input should be treated as safe or unsafe. These tasks pose significant challenges to the multimodal reasoning and safety alignment capabilities of models. In contrast, jailbreak robustness benchmarks involve adversarial attacks, such as the inclusion of maliciously crafted prompt, aimed at circumventing the safety constraints of models. Given that these two task categories examine safety alignment from different perspectives, we analyze these results separately. Further details regarding the datasets are provided in Appendix A.1.
@@
We evaluate a total of 13 MLRMs, including both proprietary models and open-source models along with their corresponding base models, such as Kimi-VL-Thinking Team et al. (2025), R1-Onevision Yang et al. (2025), Mulberry Yao et al. (2024), and LlamaV-o1 Thawakar et al. (2025). Detailed information of the models are provided in Appendix A.3. For reasoning models that have undergone extensive fine-tuning through either supervised learning or reinforcement learning, we additionally evaluate their corresponding base models. This approach allows us to trace safety-related changes across model variants. All experiments are conducted using two NVIDIA A100-80G GPUs.
@@
For jailbreak robustness benchmarks, we adopt the standard Attack Success Rate (ASR) metric, which quantifies the percentage of instances where the model produces harmful outputs with its safety mechanisms circumvented. A lower ASR indicates stronger safety alignment. For safety-awareness benchmarks, we follow established evaluation protocols to compute a safety score, where higher scores denote better safety performance. Further evaluation settings are detailed in Appendix A.1. Following the settings of MM-SafetyBench Liu et al. (2024b), MSSBench Zhou et al. (2024), and SIUO Wang et al. (2025) utilizing OpenAI API, we employ GPT-4o-mini Hurst et al. (2024) as the evaluator and adopt identical configurations to ensure consistency across assessments.
@@ 2.2 Safety Results and Inspection
Finding 1 ($\Rightarrow$ RQ1): The safety performance of MLRMs requires improvement and exhibits a notable safety degradation phenomenon.
@@
As shown in Table 1, existing multimodal reasoning models exhibit significant safety vulnerabilities, with most open-source models performing unsatisfactorily across various benchmarks. In particular, Kimi-VL-Thinking, Skywork-R1V and LLaVA-CoT successfully defend against only about 15% of malicious queries on FigStep. This finding highlights the necessity of ensuring safety alignment alongside improvements in multimodal reasoning capabilities.
Safety Degradation.
Most reasoning models exhibit a notable increase in ASR on jailbreak robustness benchmarks relative to their corresponding foundation models, with an average increase of approximately 10.57%. This suggests that the internal safety alignment mechanisms are compromised during the process of enhancing reasoning capabilities, thereby rendering the models more susceptible to jailbreak attacks. In contrast, the degradation of safety performance is less evident in safety-awareness benchmarks. In some cases, models even demonstrate improved safety performance after engaging in extended thought process. A detailed analysis of this phenomenon is presented in Finding 2.
@@
Performance on Safety-Awareness.
As shown in Table 1, the safety score of models such as Kimi-VL-Thinking Team et al. (2025), Mixed-R1 Xu et al. (2025), and SophiaVL-R1  Fan et al. (2025) are higher than their corresponding base models on safety-awareness benchmark. The effect is particularly pronounced for SophiaVL-R1 model, which achieves an 11.38% improvement on the SIUO benchmark. As illustrated in Figure 2, there is a difference in focusing center according to the responses generated by these models and their corresponding base models. Reasoning models, when engaging in long reasoning, tend to focus more on potentially unsafe details present in both the visual and textual modalities. These models improve the safety performance through internal self-reflection and error-correction mechanisms. This encouraging finding offers a promising direction for mitigating safety degradation by leveraging the innate reasoning capabilities of the models themselves.
@@
Performance of Thought Process.
Figure 5: Examples of the safety inconsistency between the thought process and the final answer. The harmful content is marked with the red font.
Finding 5 ($\Rightarrow$ RQ3): MLRMs attempt to eliminate unsafe content in the thought process via self-reflection, while there are also situations where unsafe answers are generated even if the thought process is safe.
@@
As shown in Table 3, the results reveal that the potential risk arising from the thought process does not always show up consistently in the final answer. For example, the safe thought process leads to the unsafe answer in QVQ-Preview, with respect to 46.60% proportion. As shown in Figure 5(a), in specific examples, when the model encounters unsafe intents, the reasoning process focuses merely on the directivity for answering the question, without generating explicitly harmful actions. The inconsistency between safe thought and unsafe answer is due to the potential safety risks inherent in the model. On the contrary, the unsafe thought process also results in the safe final answer. It is attributed to the internal self-reflection and error-correction mechanisms of MLRMs, which enables models to eventually recognize the harmful nature of the intended prompt as shown in Figure 5(b). The comparative results suggesting that we need to not only eliminate unsafe content in the reasoning process or answer section, but also ensure consistency between the thought process and the final answer.
@@ 3 Data Construction
Although existing studies have introduced several datasets for multimodal safety alignment, most of them Zong et al. (2024); Ding et al. (2025b) consist of brief responses that lack explicit thought process. The fine-tuning method based on these datasets struggles with retaining the inherent reasoning-chain advantages of MLRMs. Furthermore, arriving at conclusions without engaging in a safety-oriented thought process weakens the effectiveness and robustness of safety defense mechanisms. To address these issues, we propose a method for building a multimodal alignment dataset named TiS (Think in Safety), distinguished by its safety-oriented thought process and various safety-related topics. Through this method, we aim to leverage the innate reasoning capabilities of MLRMs to improve its safety alignment.
As shown in Figure 6, we employ a multi-stage pipeline to construct our safety alignment dataset TiS. We begin by collecting safety-related topics and generating image captions, then explicitly incorporating long CoT reasoning into question answering. After a filtering procedure, we finally obtain the dataset. To the best of our knowledge, TiS is the first alignment dataset with the ability to retain reasoning chain for MLRMs. We provide a detailed illustration of the entire pipeline in Appendix B.1.
@@
Due to the limitation of existing multimodal safety alignment datasets, we aim to construct a dataset enriched with safety-oriented thought process. However, considering the inherent safety vulnerabilities of existing models and the resource constraints, constructing such a dataset entirely from scratch is largely impractical. Therefore, we augment existing datasets with short responses by adding structured and safety-oriented thought process. Specifically, we select two multimodal safety alignment datasets as our original data sources: VLGuard Zong et al. (2024) and SPA-VL Zhang et al. (2024). Rather than directly utilizing the entire datasets, we retain only a subset of instances to ensure a balanced distribution across diverse safety-related topics.
@@
To obtain a reliable thought process, we utilize the text-only reasoning model Deepseek-R1 Guo et al. (2025). This is because multimodal instruction models often struggle to produce coherent reasoning, while MLRMs tend to suffer from alignment collapse under various conditions. Meanwhile, close-source proprietary models are equipped with defensive safety guardrail, which hinders access to their internal reasoning processes particularly for safety-related questions. To address the modality gap inherent in this text-only models, the visual content is converted into detailed image captions using Qwen2.5-VL-72B Bai et al. (2025).
@@
To ensure the thought process generated by LRMs is aligned with human values, we explicitly add safety guidelines in the generation process, inspired by the concept of deliberative alignment Guan et al. (2024). Previous methods often require models to infer implicit safety rules from large volumes of training examples, which suffers from low efficiency and limited generalization. By contrast, we provide clear and structured safety guidelines to directly guide the model in generating safety-oriented thought process, encouraging MLRMs to think explicitly with safety considerations when producing responses. Besides, we develop the safety guidelines customized to the specific characteristics and risk profiles of each category. More details and prompt used in this step are provided in Appendix B.3.
@@
After obtaining the safety-oriented reasoning processes, we perform an additional modification and filtering step to ensure the quality and correctness of the data, making them better suited for safety alignment. This step addresses issues such as inaccuracies (e.g., responses that explicitly state “according to the caption”) and misalignments in reasoning. We collaboratively employ Deepseek-V3 Liu et al. (2024a) alongside human annotators to assess data quality and filter inappropriate instances.
@@ 4.2 Results
As shown in Table 4, both R1-Onevision and LLaVA-CoT demonstrate improved safety alignment after fine-tuning on TiS, substantially outperforming prior datasets. Specifically, our dataset TiS enhances safety performance on FigStep and SIUO by at least 10% compared to the best alternative baseline, effectively enabling the models to leverage their reasoning capabilities for deeper analysis and unsafe intention detection. Furthermore, since our dataset incorporates thought process that closely align with the data distribution used in MLRM training, the fine-tuned model preserved the original capability to generate coherent reasoning pathways. More case studies are provided in Appendix D.4.
@@ 4.3 Discussion
To investigate whether the proposed thought process can further enhance the safety performance of MLRMs, we conduct experiments by removing the thought component from TiS dataset, retaining only the answer portion. As shown in Table 4, retaining the thought process leads to more substantial improvements in safety performance compared to using answer-only data, with the exception of MM-SafetyBench. For certain instances of MM-SafetyBench, the model’s long reasoning leads it to categorize the input as neutral, which leads to the increase of ASR.The answer-only data also exhibits favorable performance, suggesting that the construction of more comprehensive and higher-quality datasets facilitates the safe fine-tuning of MLRMs. However, this can lead to generated responses that not only lack the thought process but also consist solely of brief replies such as “I’m sorry, I can’t assist it.”.More case studies are provided in Appendix D.3. Furthermore, this also demonstrates the feasibility of leveraging the reasoning capabilities of MLRMs to enhance the alignment with safety objectives.
@@
To further assess the impact of Tis on enhancing the safety of internal reasoning, we separately evaluate the safety of both thought process and final answer, as shown in Table 5. The results demonstrate the effectiveness of TiS in safeguarding both thought and answer response, with a substantial reduction in the ASR across all categories. Notably, TiS effectively reduces the proportion of cases in which both thought process and final answer are unsafe, decreasing from 42.80% to 0.20% on R1-Onevision, and from 54.60% to 0.40% on LLaVA-CoT. Furthermore, in scenarios where unsafe content appears solely in the intermediate thinking, our fine-tuning approach also achieves a marked improvement, reducing such instances from 14.20% to 2.60%. These results indicate that TiS largely enhances the reliability and safety of the model’s internal reasoning process.